A company experiences annual demand of 2,940 units for an item that it purchases. The rate of demand per day is very stable, with very little variation from day to day. The item costs $60 when purchased in quantities less than 160 and $58 for 160 or more. Ordering costs are $50 and the carrying cost is 25 percent.

a. What will be the total costs for each alternative? (Round your intermediate calculations and final answers to the nearest dollar amount.)

b. How much should the company buy each time an order is placed?

c. How much the company can save by placing the order? (Round your answer to the nearest dollar amount.)
Answer #1

Ans:

Annual demand=D = 2940

Ordering cost=Co = $50

Carrying cost= H = 25% = 0.25

A)

When c= $60

EOQ = Sqrt((2 * D * Co)/(H*C)) = Sqrt((2 * 2940 * 50)/(0.25*60)) = Sqrt294000/15=Sqrt19600=140

It lies in the range (0 -160). Therefore EOQ = 140

Now Calculate the Total Cost for each cost level.

Total Cost = (D* C) + ((D/Q)*Co) + (Q/2)*C*H))

When C = $60, Q = EOQ = 140

TC = (2940*60) + ((2940/140)*50) + ((140/2)*0.25*60) = $ 178,500
=

When C = $58, Q = 160

TC = (2940*58) + ((2940/160)*50) + ((160/2)*0.25*58) = $ 172598.75=172599

B)

The company should buy 165 or more units

C)

The company can save ($ 178,500-172599) = $5901 annually.